You have a 1.5 m long copper wire. You want to make an N-turn current loop that generates a 2.3 mT magnetic field at the center when the current is 2.4 A . You must use the entire wire.

What will be the diameter of your coil?
Answer #1

B = uoNI/D

where N = number of turns

Length = pi*D*N

N = 1.5 / pi*D

2.3*10^-3 = 4pi*10^-7*1.5*2.4 / pi*D^2

D = 0.044 m

diameter of the coil is 4.4 cm